Suppose you have an investment with the following expected cash flows:

Years Investment A
0 -10000
1 3000
2 3000
3 6000

The IRR of this project is 8.55% per year. What you would have at the end of the third year if you rein­vested each cash flow at 8.55%?
Answer #1

IRR assumes reinvestment at IRR..8.55% is IRR so reinvesting at 8.55% means reinvesting at IRR. This would give value at the end of the third year=Initial cost*(1+IRR)^3=10000*1.0855^3=12790.56